The standard enthalpy of formation of H2O (l) is -286 kJ/mol and the standard enthalpy of combustion of ethane is -1560 kJ/mol. Given that the enthalpy change when one mole of ethene is redcued by hydrogen gas to ethane is - 138kJ/mol, calculate the standard enthalpy of combustion of ethene.
1
Expert's answer
2020-10-12T13:28:03-0400

"1) H_2(g) + \frac{1}{2}O_2(g) \rightarrow H_2O(l)" "-286 \frac{kJ}{mol}"


2) "C_2H_6(g) + \frac{7}{2} O_2(g) \rightarrow 2CO_2(g) + 3H_2O(g)" "-1560 \frac{kJ}{mol}"


3) "C_2H_4(g)+ H_2(g) \rightarrow C_2H_6(g)" "-138 \frac{kJ}{mol}"


To get the equation of etene combustion

"C_2H_4 (g) + 3O_2(g) \rightarrow 2CO_2 (g) + 2H_2O (g)"

we should "read" the first equation form right to left and then add all three euqations together

1) "H_2O(l) \rightarrow H_2(g) + \frac{1}{2}O_2(g)" "+286 \frac{kJ}{mol}"

2) "C_2H_6(g) + \frac{7}{2} O_2(g) \rightarrow 2CO_2(g) + 3H_2O(g)" "-1560 \frac{kJ}{mol}"

3) "C_2H_4(g)+ H_2(g)\rightarrow C_2H_6(g)" "-138 \frac{kJ}{mol}"


"H_2O(l) +C_2H_6(g) + \frac{7}{2} O_2(g) +C_2H_4(g)+ H_2(g) \rightarrow H_2(g) + \frac{1}{2}O_2(g)+2CO_2(g) + 3H_2O(g)+C_2H_6(g)"

"+286 -1560-138 = -1412 \frac{kJ}{mol}"


cancel the same species:

"C_2H_4(g) + 3O_2(g) \rightarrow 2CO_2(g) +2H_2O(g)" "-1412 \frac{kJ}{mol}"